OpenStudy (anonymous):

A triangle has side lengths of 34 in., 28 in., and 42 in. Is the triangle acute, obtuse, or right?

OpenStudy (anonymous):

My guess is obtuse, but I'm not sure.

Parth (parthkohli):

If \(\rm (longest~side)^2 = (one ~leg)^2 + (other~leg)^2\), then right. If \(\rm (longest~side)^2 > (one ~leg)^2 + (other~leg)^2\), then obtuse. If \(\rm (longest~side)^2 < (one ~leg)^2 + (other~leg)^2\), then acute.

OpenStudy (anonymous):

Compare a^2 + b^2 to c^2 If they're equal, it's right (Pythagorean thm) If c^2 is greater than a^2 +b^2, it's obtuse If c^2 is less than a^2 +b^2, it's acute

OpenStudy (anonymous):

OK, then I think I'm right. I'll do the math first though.

OpenStudy (anonymous):

\[28\times28=784\]\[34\times34=1156\]\[784+1156=1940\]\[\sqrt{1940}=44\]So it is obtuse.

OpenStudy (anonymous):

Thank you @ParthKohli and @bernadettegu

Parth (parthkohli):

\[42^2 = 1764\]and\[28^2 + 34^2 = 1940 \]Which one is greater?

OpenStudy (anonymous):

The second one.

Parth (parthkohli):

So look at the conditions I gave.

OpenStudy (anonymous):

So because it's longer then the longest side squared, it's obtuse.

Parth (parthkohli):

The conditions are simple... see again.

Parth (parthkohli):

\[42^2 < 28^2 + 34^2\]that's all.

OpenStudy (anonymous):

Ok I see. Thanks for actually teaching me.

Parth (parthkohli):

Hmm, the answer you get is still not right.

Parth (parthkohli):

If \(\rm (longest~side)^2 < (one ~leg)^2 + (other~leg)^2\), the acute.

OpenStudy (anonymous):

Oh, I kept reading it backwards.

Parth (parthkohli):

Ah, all right...

OpenStudy (anonymous):

So it's acute not obtuse.

Parth (parthkohli):

Right.

OpenStudy (optimusdell):

The correct answer is . Acute